# Stabilizing Thompson Sampling with Point Null Bayesian Response-Adaptive Randomization

### v1

@@ 4 Point null Bayesian RAR for binary outcomes
We now consider the setting with binary outcomes, as such outcomes frequently occur in applications of Bayesian RAR (e.g., in clinical trials or A/B testing settings). Suppose that we observe data of the form $y=\{y_{C},y_{1},\dots,y_{K},n_{C},n_{1},\dots,n_{K}\}$ where $y_{i}$ denotes the number of successes out of $n_{i}$ trials in group $i\in\{C,1,\dots,K\}$ coming from a control group (index $C$) and $K$ treatment groups. All success counts are assumed to be binomially distributed with probabilities $\theta_{C},\theta_{1},\dots,\theta_{K}$, respectively, and higher values are assumed to indicate a higher benefit (e.g., a higher probability of disease recovery). The hypotheses of interest from Section 2.2 then translate into
$H_{-}\colon\theta_{C}>\theta_{i},~i\in\{1,\dots,K\}$
$H_{0}\colon\theta_{C}=\theta_{1}=\dots=\theta_{K}$
$H_{+1}\colon\theta_{1}>\theta_{i},~i\in\{C,2,\dots,K\}$
$\vdots$
$H_{+K}\colon\theta_{K}>\theta_{i},~i\in\{C,1,\dots,K-1\}$
In the approximate normal framework from Section 3, these hypotheses could be translated into hypotheses related to log odds ratios $\psi_{i}=\log\{\theta_{i}(1-\theta_{C})\}/\{(1-\theta_{i})\theta_{C}\}$, which can be estimated with logistic regression or other methods (we will provide a comparison with this approach below). However, such normal approximations can be inaccurate for small sample sizes and/or extreme probabilities close to zero/one. It is therefore preferable to compute randomization probabilities via the exact binomial distribution.
@@
To compute posterior and randomization probabilities, it is necessary to compute the marginal likelihood of the observed data $y$ under the different hypotheses. The null hypothesis $H_{0}$ is no longer a simple hypothesis but requires specification of a prior for the common probability $\theta_{C}$. Assuming a beta prior $\theta_{C}\mid H_{0}\sim\mathrm{Beta}(a_{0},b_{0})$, the marginal likelihood of the observed data is
$\Pr(y\mid H_{0})=$ $\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{0}+\sum_{j\in\{C,1,\dots,K\}}y_{i},b_{0}+\sum_{j\in\{C,1,\dots,K\}}n_{i}-\sum_{j\in\{C,1,\dots,K\}}y_{i})}{\mathrm{B}(a_{0},b_{0})}$
@@
with $\mathrm{B}(a,b)=\int_{0}^{1}t^{a-1}(1-t)^{b-1}\mathrm{d}t$ the beta function. Under the remaining hypotheses, it is natural to assume independent beta priors $\theta_{i}\sim\mathrm{Beta}(a_{i},b_{i})$ for $i\in\{C,1,\dots,K\}$, and truncate their support to the space of the corresponding hypothesis. This leads to the marginal likelihoods
$\Pr(y\mid H_{+i})=$ $\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{j}+y_{j},b_{j}+n_{j}-y_{j})}{\mathrm{B}(a_{j},b_{j})}$
$\times\frac{Q_{i}(a_{C}+y_{C},a_{1}+y_{1},\dots,a_{K}+y_{K},b_{C}+n_{C}-y_{C},b_{1}+n_{1}-y_{1},\dots,b_{K}+n_{K}-y_{K})}{Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})}$ (9)
with
$Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$ $=\Pr(\theta_{i}=\max\{\theta_{C},\theta_{1},\dots,\theta_{K}\}\mid a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$
$=\int_{0}^{1}p(\theta_{i}\mid a_{i},b_{i})\times\prod_{j\in\{C,1,\dots,K\}\setminus\{i\}}\Pr(\theta_{j}<\theta_{i}\mid a_{j},b_{j})\,\mathrm{d}\theta_{i}$
$=\int_{0}^{1}\frac{\theta_{i}^{a_{i}-1}(1-\theta_{i})^{b_{i}-1}}{\mathrm{B}(a_{i},b_{i})}\times\prod_{j\in\{C,1,\dots,K\}\setminus\{i\}}I_{\theta_{i}}(a_{j},b_{j})\,\mathrm{d}\theta_{i},$
where $I_{x}(a,b)=\{\int_{0}^{x}t^{a-1}(1-t)^{b-1}\,\mathrm{d}t\}/\mathrm{B}(a,b)$ is the regularized incomplete beta function, also known as the cumulative distribution function of the beta distribution. These can be efficiently computed with numerical integration and standard implementations of the regularized incomplete beta function (e.g., stats::pbeta in R). The marginal likelihood of the data under $H_{-}$ can similary be obtained from (9) with $i=C$.
For a specified prior probability of the null hypotheses $\Pr(H_{0})$, we may again distribute the remaining prior probability among the other hypotheses by
$\Pr(H_{+i})=\{1-\Pr(H_{0})\}\{1-Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})\}$
to ensure that the averaged prior is again the beta prior that was truncated in the first place. Plugging these marginal likelihoods and prior probabilities into equation (1) produces posterior probabilities, which in turn can be used for obtaining randomization probabilities. Similarly, ratios of marginal likelihoods can be taken to obtain Bayes factors for monitoring accumulating evidence.
Figure 6: Evolution of Bayesian RAR probabilities for 3 treatments and a control group. In each step, one observation from each group is simulated from a binomial distribution with probability as indicated in the panel titles (treatment 1 is the most effective). Randomization probabilities are computed assuming binomial likelihoods with independent uniform priors or a normal approximation to the vector of log odds ratios obtained from logistic regression along with a multivariate normal prior with variances 11 and correlations $0.5$.
Figure 6 illustrates RAR probabilities for simulated binomial data with a control and $K=3$ treatment groups. Randomization probabilities were computed assuming exact binomial likelihoods with uniform priors for the probabilities or a normal approximation to the vector of log odds ratios obtained from logistic regression along with a multivariate normal prior. We can see that the exact (solid lines) and approximate (dashed lines) probabilities are not too far from each other in most cases even though the true probabilities from which the data are simulated are rather small. The RAR probabilities with $\Pr(H_{0})=0.5$ remain considerably closer to 25% (equal randomization) compared to the RAR probabilities with $\Pr(H_{0})=0$ (Thompson sampling), which shows the largest variability. Since the true probability in treatment group 1 is the highest, all randomization probabilities converge towards $100\%$ for treatment 1, as expected, while they decrease to $0\%$ for the other treatments.
@@ 5 Reanalyzing the ECMO trial
The ECMO trial (Bartlett et al.,, 1985) investigated the efficacy of ECMO (extracorporeal membrane oxygenation) treatment in the critical care of newborns. It was the first prominent clinical trial to use a “randomized play-the-winner” (RPW) RAR design (Wei and Durham,, 1978). While some earlier non-randomized studies had shown a substantial treatment effect, a randomized study was required to confirm this. However, this presented an ethical dilemma, as the investigators were convinced that the risk of death would be much higher in the control group than in the ECMO group. To mitigate this, the RPW design was chosen.
@@
The outcome of the trial was extreme: The first newborn was randomized to receive ECMO treatment and survived. The second newborn was randomized to receive the control treatment and died. The ten subsequently enrolled newborns were all randomized to ECMO, and all survived (Bartlett et al.,, 1985). The trial was then stopped for efficacy and its results published. However, the unusual outcome sparked intense debates about whether the trial was even a proper randomized clinical trial. Several subsequent trials were conducted, all re-establishing the effectiveness of ECMO, which is now a standard treatment (Bartlett,, 2024). In the following, we will reanalyze the ECMO data.
@@
Figure 7: Evolution of Bayesian RAR randomization probabilities (top plot) and posterior probability of a beneficial ECMO treatment effect (bottom plot) for data from the ECMO trial.
@@
The top plot in Figure 7 shows Bayesian RAR probabilities computed from the ECMO data sequence. The normal approximation and the exact binomial methods were used, as well as the RPW method that was originally used in the ECMO trial. A standard normal prior was assigned to the log odds ratio for the normal method, while uniform priors were assigned to the probabilities for the binomial method. For the normal method, log odds ratio estimates were computed with a Yates’ correction (i.e., adding a half to each cell) to avoid issues with zero cells. We can see that this correction induces a slight anomaly at the start of the study as the probabilities from the normal method slightly decrease after the first observation despite the first patient surviving the ECMO treatment. This does not happen for the exact method whose randomization probability increase after every additional patient, which only happens for the normal approximation after the second patient. Furthermore, the randomization probabilities differ notably between the normal and exact methods, even at the end of the study. This presumably happens because the influence of the differing priors remains relatively high after observing only 12 patients.
@@
Comparing the randomization probabilities for different prior probabilities of the null hypothesis for each method, we can see that $\Pr(H_{0})=0$ (Thompson sampling) shows the most extreme randomization probabilities that rapidly increase to 100%, whereas $\Pr(H_{0})=1$ (equal randomization) leaves the probabilities completely static at 50%. In between, the randomization probabilities gradually shift from 50% to the Thompson sampling probabilities. The RPW probabilities (black squares) are relatively close to the probabilities from the exact method with $\Pr(H_{0})=0.5$ at the beginning of the study and become closer to the normal approximate method with $\Pr(H_{0})=0.75$ at later time points.
@@
The bottom plot in Figure 7 shows the corresponding posterior probability of the ECMO treatment being effective. Depending on the prior probability $\Pr(H_{0})$, we can see that the posterior probability at the end of the trial may be very high (e.g., $\Pr(H_{+}\mid y)=0.99$ for $\Pr(H_{0})=0$), or only moderately favoring ECMO (e.g., $\Pr(H_{+}\mid y)=0.86$ for $\Pr(H_{0})=0.5$). From a Bayesian perspective, stopping the trial seems thus only a sensible decision if the prior probability of equal effects was low. Conversely, more evidence would have been needed if the prior probability had been higher, for example, if the prior probability were $\Pr(H_{0})=0.5$, representing a priori equipoise.
@@ 6 Simulation study
We conducted a simulation study to evaluate the performance of point null Bayesian RAR for different values of the prior probability of the null hypothesis $\Pr(H_{0})$, and compare it to Thompson sampling (potentially modified with burn-in periods, probability capping, power transformations) and equal randomization. Considered patient performance measures were the rate of successes, the rate of extreme randomization probabilities, and sample size imbalance. Additionally, bias and coverage were considered to evaluate the performance of rate difference point estimates and confidence intervals under RAR, while the type I error rate and power of the corresponding tests were used to quantify hypothesis testing performance under RAR. A binomial data-generating mechanism was used which was partially based on the simulation studies from Robertson et al., (2023), Thall and Wathen, (2007), and Wathen and Thall, (2017). Detailed description of the design and results of the simulation study following the structured ADEMP approach (Morris et al.,, 2019; Siepe et al.,, 2024) are provided in Appendix B. A supplemental website provides an interactively explorable results dashboard (https://samch93.github.io/brar/).
Across all simulations, we observed a trade-off between patient benefit and parameter estimation / hypothesis testing operating characteristics. Some methods performed better in terms of patient benefit but had worse bias, coverage, type I error rate, and power while others performed better in terms of bias, coverage, type I error rate, and power but had worse patient benefit. This trade-off is well described in the RAR literature (Hu and Rosenberger,, 2003; Zhang and Rosenberger,, 2005; Williamson and Villar,, 2019; Robertson et al.,, 2023).
@@
The main result of the simulation study regarding the newly proposed method was that, under most conditions, point null Bayesian RAR with a high prior probability of the null hypothesis $\Pr(H_{0})=0.75$ showed similar operating characteristics to Thompson sampling with capped randomization probabilities at 10% and 90% and a power transformation with $c=i/(2n)$, where $i$ is the current sample size and $n$ is the maximum sample size. Both point null Bayesian RAR with a prior probability of $\Pr(H_{0})=0.75$ and modified Thompson sampling could mitigate some issues with ordinary Thompson sampling. For instance, they exhibited less negative sample size imbalance, biased parameter estimates, undercoverage, and inflated type I error rates. However, this improvement came at the cost of worse patient benefit performance compared to unmodified Thompson sampling. For instance, the mean success rate was lower, though still considerably better than equal randomization in most cases. Most importantly, the variability of randomization probababilitis was much reduced with $\Pr(H_{0})=0.75$, producing rarely negative imbalances (i.e., a large proportion of patients randomized to an inferior treatment). Setting a lower but positive prior probability than $\Pr(H_{0})=0.75$ produced operating characteristics comparable to those from less extreme modifications of Thompson sampling, such as a power transformation with $c=1/2$. In sum, the simulation study demonstrated that point null Bayesian RAR has comparable statistical properties to Thompson sampling with common ad hoc modifications.

### v2

@@ 4 Null hypothesis Bayesian RAR for binary outcomes
Suppose now that we observe binary data of the form $y=\{y_{C},y_{1},\dots,y_{K},n_{C},n_{1},\dots,n_{K}\}$ where $y_{i}$ denotes the number of successes out of $n_{i}$ trials in group $i\in\{C,1,\dots,K\}$ coming from a control group ($C$) and $K$ treatment groups. All success counts are assumed to be binomially distributed with probabilities $\theta_{C},\theta_{1},\dots,\theta_{K}$, respectively, with higher values indicating greater benefit. The hypotheses from Section 2.1 then translate into $H_{-}\colon\theta_{C}=\max\{\theta_{C},\theta_{1},\dots,\theta_{K}\}$ vs. $H_{0}\colon\theta_{C}=\theta_{1}=\dots=\theta_{K}$ vs. $H_{+i}\colon\theta_{i}=\max\{\theta_{C},\theta_{1},\dots,\theta_{K}\}$ for $i=1,\dots,K$. In the normal framework from Section 3, these could be translated into hypotheses related to log odds ratios, which can be estimated with logistic regression or other methods. However, normal approximations can be inaccurate for small sample sizes or extreme probabilities, so exact binomial computation is preferable.
@@
The null hypothesis $H_{0}$ requires specification of a prior for the common probability $\theta_{C}$. Assuming a beta prior $\theta_{C}\,|\,H_{0}\sim\mathrm{Beta}(a_{0},b_{0})$, the marginal likelihood under $H_{0}$ is
$\Pr(y\,|\,H_{0})=$ $\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{0}+\sum_{i\in\{C,1,\dots,K\}}y_{i},b_{0}+\sum_{i\in\{C,1,\dots,K\}}n_{i}-y_{i})}{\mathrm{B}(a_{0},b_{0})}$
@@
with beta function $\mathrm{B}(\cdot,\cdot)$. Under the other hypotheses, it is natural to assign independent beta priors $\theta_{i}\sim\mathrm{Beta}(a_{i},b_{i})$ for $i\in\{C,1,\dots,K\}$, and truncate their support to the space of the corresponding hypothesis. This leads to the marginal likelihood of the data
$\Pr(y\,|\,H_{i})$ $=\prod_{j\in\{C,1,\dots,K\}}\binom{n_{j}}{y_{j}}\times\frac{\mathrm{B}(a_{j}+y_{j},b_{j}+n_{j}-y_{j})}{\mathrm{B}(a_{j},b_{j})}$
$\times\frac{Q_{i}(a_{C}+y_{C},a_{1}+y_{1},\dots,a_{K}+y_{K},b_{C}+n_{C}-y_{C},b_{1}+n_{1}-y_{1},\dots,b_{K}+n_{K}-y_{K})}{Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})}$
under $H_{i}\in\{H_{C},H_{+1},\dots,H_{+K}\}$ and with
$Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$ $=\int_{0}^{1}\frac{\theta_{i}^{a_{i}-1}(1-\theta_{i})^{b_{i}-1}}{\mathrm{B}(a_{i},b_{i})}\times\prod_{j\in\{C,1,\dots,K\}\setminus\{i\}}I_{\theta_{i}}(a_{j},b_{j})\,\mathrm{d}\theta_{i},$
where $I_{x}(a,b)$ is the CDF of the beta distribution. These marginal likelihoods can be efficiently computed with numerical integration. For priors with integer hyperparameters, there even exist closed-form solutions (Howard,, 1998; Miller,, 2015).
For a specified $\Pr(H_{0})$, we may again distribute the remaining prior probability among the other hypotheses by $\Pr(H_{+i})=\{1-\Pr(H_{0})\}\times Q_{i}(a_{C},a_{1},\dots,a_{K},b_{C},b_{1},\dots,b_{K})$ to ensure that the averaged prior is again the beta prior. Specifying uniform priors ($a_{j}=b_{j}=1~\forall~j\in\{C,1,\dots,K\}$) is an intuitive default that ensures that each treatment receives equal prior probability. Plugging the marginal likelihoods and prior probabilities into (1) produces posterior probabilities, which in turn can be used to obtain randomization probabilities.
@@ 5 Reanalyzing the ECMO trial
The ECMO trial (Bartlett et al.,, 1985; Bartlett,, 2024) investigated ECMO (extracorporeal membrane oxygenation) treatment in critically ill newborns. While earlier non-randomized studies suggested a substantial treatment effect, a randomized study was required to confirm this. However, investigators were convinced the mortality risk was much higher in the control group, posing an ethical dilemma. To mitigate this, a “randomized play-the-winner” (RPW) RAR design (Wei and Durham,, 1978) design was chosen.
@@
The trial’s outcome was extreme: The first newborn was randomized to ECMO and survived. The second was randomized to the control and died. All ten subsequently enrolled newborns were randomized to ECMO and survived. The trial was then stopped for efficacy, yet its unusual outcome sparked debate about whether it constituted a proper randomized trial. Later trials confirmed the efficacy of ECMO, which is now a standard treatment.
@@
Figure 5: Evolution of Bayesian RAR randomization probabilities (a) and posterior probability of a beneficial ECMO treatment effect (b) for data from the ECMO trial. Plot (c) shows the probability of observing the ECMO allocation sequence as a function of $\Pr(H_{0})$.
@@
Figure 5a shows RAR probabilities computed from the ECMO data. The normal approximation and exact binomial methods were used, as well as the RPW method originally used in the ECMO trial. A standard normal prior was assigned to the log odds ratio for the normal method, while uniform priors were assigned to the probabilities for the binomial method. Log odds ratio estimates were computed by adding a half to each cell to avoid zero cell issues. This correction induces a slight anomaly, as the probabilities from the normal method slightly decrease after the first newborn despite that it surived. This does not happen for the exact method whose probabilities increase after every additional patient. Furthermore, the randomization probabilities differ notably between the normal and exact methods, even at the end of the study. This presumably happens because the influence of the differing priors remains relatively high after observing only 12 patients.
@@
Comparing the randomization probabilities for different prior probabilities of the null hypothesis, $\Pr(H_{0})=0$ (Thompson sampling) shows the most extreme randomization probabilities that rapidly increase to 100%, whereas $\Pr(H_{0})=1$ (equal randomization) leaves the probabilities completely static at 50%. In between, the probabilities gradually shift from 50% to the Thompson sampling probabilities. The RPW probabilities (black squares) are relatively close to the exact method with $\Pr(H_{0})=0.5$ at the beginning of the study and become closer to the normal method with $\Pr(H_{0})=0.75$ at later times.
@@
Figure 5b shows the corresponding posterior probability of the ECMO treatment being effective. Depending on $\Pr(H_{0})$, the posterior probability at the end of the trial may be very high (e.g., $\Pr(H_{+}\,|\,y)=0.99$ for $\Pr(H_{0})=0$), or only moderately favoring ECMO (e.g., $\Pr(H_{+}\,|\,y)=0.86$ for $\Pr(H_{0})=0.5$). Stopping the trial seems thus only a sensible decision if the prior probability of no ECMO effect was low.
Figure 5c shows the probability of observing the ECMO allocation sequence under Bayesian RAR with different $\Pr(H_{0})$. The probability is highest for $\Pr(H_{0})=0$ (Thompson sampling) and decreases with increasing $\Pr(H_{0})$ until $(1/2)^{12}=0.024\%$ for $\Pr(H_{0})=1$ (equal randomization). Interestingly, the probability under RPW is also rather low (2.8%). Hence, observing the ECMO allocation sequence under Bayesian RAR would only be less likely than under RPW for $\Pr(H_{0})>0.6$ (exact method) and $\Pr(H_{0})>0.55$ (normal method), respectively.
@@ 6 Simulation study
We conducted a simulation study to evaluate the performance of null hypothesis Bayesian RAR for different values $\Pr(H_{0})$, and compare it to Thompson sampling (potentially modified with burn-in periods, probability capping, power transformations), equal randomization, the Gittins index (Gittins et al.,, 2011), and Bayes upper confidence bound (UCB) (Kaufmann et al.,, 2012). Considered patient benefit measures were the mean rate of successes, the rate of extreme randomization probabilities, and sample size imbalance in favor of the inferior treatment. Bias and coverage were used to evaluate performance of response rate difference point estimates and confidence intervals under RAR, while the type I error rate and power of the corresponding tests were used to quantify hypothesis testing performance. A binomial data-generating mechanism was used. The response probability in the control group was fixed to $\theta_{C}=0.25$ and varied in the first treatment group $\theta_{1}\in\{0.25,0.35,0.45\}$ to assess the behavior of the methods for different effect sizes $\text{RD}_{1}=\theta_{1}-\theta_{C}$. In conditions with two or three treatment groups, the corresponding probabilities were set to $\theta_{2}=\theta_{3}=0.3$. More detailed descriptions of the design and results of the simulation study are provided in the supplement and an online results dashboard (https://samch93.github.io/brar/).
A trade-off between patient benefit and inferential operating characteristics was observed across most settings. Methods that improved patient benefit through more aggressive allocation (Gittins index, Bayes UCB, and Thompson sampling) often had worse bias, coverage, type I error rate, and power, whereas more conservative methods tended to preserve inferential validity at the cost of reduced patient benefit. Figure 6 illustrates this trade-off for the mean success rate and coverage. This trade-off is well documented (Hu and Rosenberger,, 2003; Zhang and Rosenberger,, 2005; Williamson and Villar,, 2019; Robertson et al.,, 2023). However, there were also settings with a small “window” in which patient benefit could vary while inferential performance remained essentially unchanged. For example, with a burn-in of 50, two treatment groups ($K=2$), and a large effect size ($\text{RD}_{1}=0.2$), coverage was roughly constant across $\Pr(H_{0})$, whereas the mean success rate changed noticeably. Increasing $\Pr(H_{0})$ reduced imbalance toward the inferior treatment, though negative imbalance was also strongly influenced by effect size. In conditions with large effect size ($\text{RD}_{1}=0.2$), the proportion of negative imbalance was often lower for RAR methods (including Thompson sampling) than equal randomization, thus confirming the results from Robertson et al., (2023). These findings highlight the importance of evaluating operating characteristics during trial design, since appropriately specified RAR designs may enable improved patient benefit without substantial loss of inferential performance.
Figure 6: Plot (a) shows the mean rate of successes (i.e., the number of successes in a study divided by its sample size averaged over all 10’000 simulation repetitions). The maximum Monte Carlo standard error (MCSE) is 0.051%. Plot (b) shows the empirical coverage rate of the 95% Wald confidence interval for the response rate difference $\text{RD}_{1}$. The maximum MCSE is 0.46%.
@@
Under most conditions, Bayesian RAR with $\Pr(H_{0})=0.75$ showed similar operating characteristics to Thompson sampling with capped randomization probabilities at 10%/90% and power transformation $c=i/(2n)$, where $i$ is the current sample size and $n$ is the maximum sample size. Both mitigated issues with ordinary Thompson sampling, exhibiting less negative sample size imbalance, less bias, improved coverage, and reduced type I error rates. This came at the cost of worse patient benefit performance, for instance, lower mean success rate, though still better than equal randomization in most cases. Smaller values of $\Pr(H_{0})$ produced operating characteristics comparable to less extreme modifications, such as $c=1/2$. Finally, the results were similar for exact binomial and approximate normal versions of the method (supplement), suggesting some robustness to prior and model choice.
While these results demonstrated that null hypothesis Bayesian RAR has comparable properties to capped Thompson sampling for realistic trial sizes, it is important to note that the two are expected to differ asymptotically. Capped Thompson sampling enforces bounds on randomization probabilities, whereas uncapped Bayesian RAR probabilities can converge to 100% or 0%. The observed similarity is thus only a finite-sample phenomenon.